OpenStudy (moonlight93):

Solve the triangle, given the following : angle A= 30, side b= 6 , and side c=8 find angle C, Angle B, and side a?

OpenStudy (campbell_st):

well the triangle looks like  find a by using the law of cosines then find B using the law of Sines then find C using angle sum of a triangle hope it helps

OpenStudy (moonlight93):

a=9.2 , i tried to use law of sine , but the answer didn't make sense

OpenStudy (moonlight93):

sinB/6 = sine30/9.23 ..... couldn't get a realistic solution from it

OpenStudy (campbell_st):

the law of sines says \[\frac{\sin(30)}{9.2} = \frac{\sin(B)}{6}\]

OpenStudy (campbell_st):

so \[\sin(B) = \frac{6 \times \sin(30)}{9.2}\] you will get a decimal less than 1... then its \[B = \sin^{-1}(answer) \]

OpenStudy (campbell_st):

and you may like to check your solution to a

OpenStudy (moonlight93):

I got a negative solution, won't that be undefined .. because the limit is 0-1?

OpenStudy (campbell_st):

ok.... go back to the law of cosines \[a^2 = 6^2 + 8^2 - 2 \times 6 \times 8 \times \cos(30)\]

OpenStudy (campbell_st):

by my calculations its not 9.2

OpenStudy (campbell_st):

if you look at the triangle the angle of 30 is opposite the longest side 9.2 so B and C would need to be less than 30... so its not a triangle

OpenStudy (campbell_st):

still there @moonlight93

OpenStudy (moonlight93):

so would that be "can not be solved" ?

OpenStudy (campbell_st):

no... go back and redo the law of cosines... as your answer is incorrect for a I posted it correctly above...

OpenStudy (moonlight93):

I still don't get it. i've been trying over and over. do you know the answer of angle B ?

OpenStudy (campbell_st):

ok... \[a^2 = 6^2 + 8^2 - 2 \times 6 \times 8 \times \cos(30)\] \[a^2 = 100 - 83.14\] what is the value of a...?

OpenStudy (moonlight93):

a=4.106

OpenStudy (moonlight93):

but why did you get a 83.14 for the right side? I got 100-14.808 ... and that's how I got the 9.23

OpenStudy (campbell_st):

great... so now the measurements make sense then \[\frac{\sin(30)}{4.1} = \frac{\sin(B)}{6}\] so sin(B) = 0.731717

OpenStudy (campbell_st):

so then \[B = \sin^{-1}(0.731707)\] then find C using angle sum of a triangle.

OpenStudy (campbell_st):

83.2....= 2 x 6 x 8 cos(30)

OpenStudy (campbell_st):

I have to go... Hope its helped.

OpenStudy (moonlight93):

ok, thank you

OpenStudy (anonymous):

@@ is your triangle even a right triangle?

OpenStudy (moonlight93):

@campbell_st ... sorry yes it will equal 83.2 .. my calculator was in radians mode, I got that answer when i swiched to degree mode. thank you for helping :)

OpenStudy (moonlight93):

@BabiiMuffin , it doesn't say it is. we've been studying both non-right and right triangle.. though we suppose to solve as if it's a non-triangle.